Pete travels 60 miles in an hour. The first 20 miles he travels at a speed of 50 mph. Another third of the way he travels 60 mph. How fast did he travel the last 20 miles to make it to the train station in one hour. Show all work if possible.
Answer by richwmiller(17219)

r*t=d
t=d/r
20/50+20/60+20/x=1
2/5+1/3+20/x=1
6/15+5/15+20/x=1
11/15+20/x=1
20/x=4/15
300=4x
x=75 mph
check
20/50+20/60+20/75=1
6/15+5/15+4/15=15/15
6+5+4=15
ok